If f(x)=x^2-3 & g(x)=x+5, which is equivalent form of f(g(x))?
A.x^2+10x+22
B.x^3+5x^2-3x-15
C.x^2+2
D.x^2+x+2
Can someone plz answer.

Answer by jim_thompson5910(35256)

f(x)=x^2-3

f(g(x))=(g(x))^2-3

f(g(x))=(x+5)^2-3

f(g(x)) = x^2+10x+25 -3

f(g(x)) = x^2+10x+22